3. (4pts) Rewrite the series as a series with general term r (that is, of the form .x). +] (#);
Answer #1

Solution r xn-4 n +1 n=2 3n = nti h=2 303n-2 n=2 39 How, replace n by n.42 - nti xn-4 +7) +1 n +2 -2 m2 3n ht2=2 29t2 n +3 no